### v1

@@
[…]
Dan Wang Department of Electronic and Computer Engineering, The Hong Kong University of Science and Technology, Clear Water Bay, Kowloon, Hong Kong, China ().    Wei Chen22footnotemark: 2    Sei Zhen Khong Department of Electrical and Electronic Engineering, The University of Hong Kong, Pokfulam, Hong Kong, China ().    Li Qiu22footnotemark: 2
?abstractname?
In this paper, we define the phases of a complex cramped matrix to be the phases of the eigenvalues of the unitary part of an associated sectoral decomposition. While a cramped matrix admits distinct sectoral decompositions, its phases defined in this manner are unique. In particular, we show that specific forms of sectoral decompositions are indeed unique, namely the symmetric polar decomposition and generalized Cholesky factorization. Various properties of matrix phases are studied, including those of compressions, Schur complements, matrix products, and sums. Importantly, by exploiting a notion known as the compound numerical range, we establish a majorization relation between the phases of the eigenvalues of a matrix product and the phases of the individual matrices. This is then applied to investigate the rank of the sum of a matrix product and the identity matrix under phasic perturbations. The phases of the Kronecker and Hadamard products are also discussed.
@@
phases, arguments, cramped matrices, interlacing, majorization, rank robustness
{AMS}
15A18, 15A23, 15A42, 15A60, 15B48, 15B57
@@ 1 Introduction
To be more specific, the phase $\phi$ takes values in a half open $2\pi$-interval, typically $0,2\pi)$ or $(-\pi,\pi]$. If $c=0$, the magnitude $|c|$ is defined as 0, but the phase $\angle c$ is undefined.
@@
(1) $|ab|$ $=|a||b|$ (multiplicativity)
(2) $|a+b|$ $\leq|a|+|b|$ (subadditivity)
(3) $\angle(ab)$ $=\angle a+\angle b\ \ \ \mbox{ mod $2\pi$}$
(4) $\angle(a+b)$ $\in\mathrm{Co}\{\angle a,\angle b\}\ \ \ \mbox{if $|\angle a-\angle b|<\pi$},$
@@
It is well understood that an $n\times n$ complex matrix $C$ has $n$ magnitudes, served by the $n$ singular values
@@
The singular values can be obtained from a singular value decomposition $C=USV^{*}$, where $U,V$ are unitary, and $S=\mathrm{diag}\left\{\sigma_{1}(C),\sigma_{2}(C),\dots,\sigma_{n}(C)\right\}$ [12, Theorem 2.6.3]. They can also be derived from the following maximin and minimax expressions [3, Corollary III.1.2][12, Theorem 4.2.6]:
@@
Note that log-majorization is stronger than weak-majorization, i.e., $x\prec_{\log}y$ implies $x\prec_{w}y$ [16, Chapter 5, A.2.b]. For more details on the theory of majorization, we refer to [16] and the references therein. Denote the vector of eigenvalues of $C$ by
@@
It is well known that the magnitudes of the eigenvalues are bounded by the singular values in the following manner [16, Theorem 9.E.1]:
@@
It is also of interest to relate the singular values of $C$ with those of any $(n-k)\times(n-k)$ submatrix of $C$. In this regard, we have the following interlacing property [3, Corollary III.1.5]:
Let $C\in\mathbb{C}^{n\times n}$ be partitioned as $C=\begin{bmatrix}C_{11}&C_{12}\\ C_{21}&C_{22}\end{bmatrix}$, where $C_{11}\in\mathbb{C}^{k\times k}$. If $C_{11}$ is nonsingular, then the Schur complement of $C_{11}$ in $C$, denoted by $C/\penalty 50_{11}$, exists and is given by $C/\penalty 50_{11}=C_{22}-C_{21}C_{11}^{-1}C_{12}$. It can be inferred from [22, Corollary 2.3] that $\sigma(C/\penalty 50_{11})$ and $\sigma(C)$ satisfy the following interlacing relation:
The singular values of $AB$ and $A+B$ are related to those of $A$ and $B$ through the following majorization type inequalities resembling (1) and (2), respectively, [16, Theorem 9.H.1, 9.G.1.d]
(5) $\sigma(AB)$ $\prec_{\log}\sigma(A)\odot\sigma(B),$
$\sigma(A+B)$ $\prec_{w}\sigma(A)+\sigma(B),$
@@
In contrast to the magnitudes of a complex matrix $C$, there does not exist a universally accepted definition of phases of $C$. What properties should the phases satisfy? In this paper, we advocate a definition of matrix phases and derive some of their properties as counterparts to those of the singular values described above.
An early attempt [18] defined the phases of $C$ as the phases of the eigenvalues of the unitary part of its polar decomposition, as motivated by the seeming generalization of the polar form of a scalar to the polar decomposition of a matrix. As in the scalar case, ambiguity arises when defining phases for a singular $C$. Hence, only nonsingular matrices are relevant. More precisely, let the left and right polar decompositions of a nonsingular matrix $C$ be given by $C=PU=UQ$, where $P$ and $Q$ are positive definite and $U$ is unitary [12, Theorem 7.3.1]. The authors of [18] proposed to define the phases of $C$ as
(6) $\psi(C)=\angle\lambda(U).$
@@
In this paper, we shall revisit the canonical angles introduced in [9] and propose to adopt them as the phases of a complex matrix. At present, we limit the definition to the class of cramped matrices, i.e., matrices whose numerical ranges do not contain the origin.
@@ 2 Definition of phases
First, we review the concepts of numerical range, angular numerical range, and canonical angles of a matrix [13, Chapter 1]. The numerical range, also called the field of values, of a matrix $C\in\mathbb{C}^{n\times n}$ is defined as
$W(C)=\{x^{*}Cx:x\in\mathbb{C}^{n}\mbox{ with }\|x\|=1\},$
@@
If $0\notin W(C)$, then $W(C)$ is completely contained in an open half plane by its convexity. In this case, $C$ is said to be a cramped matrix. We note that [2] defined a cramped matrix to mean $0\notin W(C)$, but only when $C$ is unitary. Here we say a matrix $C$ is cramped if $0\notin W(C)$, regardless of whether it is unitary or not. Early studies of cramped unitary matrices can also be found in [20, 15], among others.
Let $C$ be a cramped matrix. It is well known [11, 5, 7] that $C$ is congruent to a unitary matrix. In particular, the unitary matrix can be chosen to be diagonal, i.e., there exist a nonsingular matrix $T$ and a diagonal unitary matrix $D$ such that
(7) $C=T^{*}DT.$
The factorization (7) is called sectoral decomposition [23]. It is clear that $W^{\prime}(D)=W^{\prime}(C)$ and thus $D$ is also cramped. Let $\delta(C)$ be the field angle of $C$, i.e., the angle subtended by the two supporting rays of $W^{\prime}(C)$ at the origin. We define the phases of $C$, denoted by $\phi_{1}(C),\phi_{2}(C),\dots,\phi_{n}(C)$, to be the phases of the eigenvalues (i.e., diagonal elements) of $D$, taking values in an interval $(\theta,\theta+\pi)$ for some $\theta\in[-\pi,\delta(C))$. The phases defined in this way coincide with the canonical angles of $C$ introduced in [9].
A sectoral decomposition for a cramped matrix is not unique. Nevertheless, the diagonal unitary matrix $D$ is unique up to a permutation, as pointed out in [23]. As such, the phases of a cramped matrix are unique. We adopt the convention of labeling the phases in such a way that
@@
Define $\phi(C)=[\phi_{1}(C)\ \ \phi_{2}(C)\ \ \dots\ \ \phi_{n}(C)]$. The uniqueness of the phases also follow directly from the maximin and minimax expressions [11, Lemma 8]:
(8) $\begin{split}\phi_{i}(C)&=\max_{\mathcal{M}:\mathrm{dim}\mathcal{M}=i}\min_{x\in\mathcal{M},\|x\|=1}\angle x^{*}Cx\\ &=\min_{\mathcal{N}:\mathrm{dim}\mathcal{N}=n-i+1}\max_{x\in\mathcal{N},\|x\|=1}\angle x^{*}Cx.\end{split}$
@@
from which we can observe that $\overline{\phi}(C)$ and $\underline{\phi}(C)$ enjoy good geometric interpretations. Consider the two supporting rays of $W(C)$. Since $0\notin W(C)$, both the supporting rays can be determined uniquely. Figure 1 illustrates an example of $W(C)$ and its supporting rays. The two angles from the positive real axis to the supporting rays are $\overline{\phi}(C)$ and $\underline{\phi}(C)$ respectively. The other phases of $C$ lie in between.
?figurename? 1: Geometric interpretation of $\overline{\phi}(C)$ and $\underline{\phi}(C)$.
It is worth noting that the class of unitoids was introduced in [14] to consist of matrices that are congruent to unitary matrices. Clearly, cramped matrices constitute a special type of unitoid matrices. A noncramped unitoid matrix admits a factorization of the form (7). However, in this case, the eigenvalues of $D$ lie on an arc of the unit circle of length no less than $\pi$. In this paper, we do not define the phases of such matrices.
{eg}
@@
is not cramped. We do not define the phases of this matrix, since in this endeavor there would clearly be an ambiguity in deciding whether the phases should be taken as $2\pi/3,0,-2\pi/3$ or $4\pi/3,2\pi/3,0$.
@@
    The phases of a cramped normal matrix are the phases of its eigenvalues.
@@
    The phases of $C$ belong to $(-\pi/2,\pi/2)$ if and only if $C$ has a positive definite Hermitian part, i.e., $(C+C^{*})/2>0$. Such matrices are said to be strongly positive real in the engineering literature [1, p. 51].
@@
(9) $\psi\left(\mbox{$\left[\begin{array}[]{cc}\cos\alpha&0\\ 0&\sin\alpha\end{array}\right]$}\mbox{$\left[\begin{array}[]{cc}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{array}\right]$}\mbox{$\left[\begin{array}[]{cc}\cos\alpha&0\\ 0&\sin\alpha\end{array}\right]$}\right)\\ =\begin{bmatrix}\arccos\displaystyle\frac{\cos\theta}{\sqrt{1-\sin^{2}\theta\cos^{2}2\alpha}}&\;-\arccos\displaystyle\frac{\cos\theta}{\sqrt{1-\sin^{2}\theta\cos^{2}2\alpha}}\end{bmatrix}$,
@@
which is negative for some $\theta$ close to $\pi/2$ and $\alpha$ away from $\pi/4$, such as $\theta=\pi/4$ and $\alpha=\pi/12$. That means the matrix is not positive real for these values of $\theta$ and $\alpha$.
The following lemma shows that the phases of a cramped matrix provide a bound on the phases of its eigenvalues, just as the gains of a matrix provide a bound on the magnitudes of its eigenvalues. It has been proved implicitly in [11, Lemma 9] and also shown in [9, Theorem 1].
Lemma 2.1.
Let $C$ be cramped with phases in $(\theta,\theta+\pi)$, where $\theta\in[-\pi,\delta(C))$. Then the phases of the eigenvalues of $C$ can be chosen so that
@@
in Example 2, where $\alpha,\theta\in(0,\frac{\pi}{2})$. One can see from the discussion therein that $\psi(C)\prec\phi(C)$. In fact, this is generally true for any cramped matrix $C$ with phases in $(\theta,\theta+\pi)$, if $\psi(C)$ takes values in the same interval, which can be inferred from [11, Lemma 9].
3 Cramped matrix decompositions
As discussed in the previous section, a cramped matrix $C$ admits a sectoral decomposition as in (7), which is not unique. If the unitary matrix in (7) is not restricted to be diagonal, one gains more flexibility. In fact, letting $T=VP$ be the unique right polar decomposition and $U=V^{*}DV$ leads to a new decomposition
@@ 3 Cramped matrix decompositions
where $P$ is positive definite and $U$ is unitary. This is called the symmetric polar decomposition (SPD). Meanwhile, letting $T=QR$ be the unique QR factorization [12, Theorem 2.1.14] with $Q$ being unitary and $R$ being upper triangular with positive diagonal elements and letting $W=Q^{*}DQ$ leads to another decomposition
@@
Theorem 3.1.
The SPD and GCF of a cramped matrix $C$ are unique.
Proof 3.2.
@@
Since the phases of a cramped matrix are unique, we can find unitary matrices $W_{1}$ and $W_{2}$ such that $U_{1}=W^{*}_{1}DW_{1}$ and $U_{2}=W^{*}_{2}DW_{2}$, where
@@
(11) $\begin{split}e^{i\phi_{[k]}}I_{n_{k}}&=\sum_{l=1}^{m}e^{i\phi_{[l]}}T^{*}_{lk}T_{lk},\quad k=1,2,\dots,m\quad\text{and}\\ 0&=\sum_{l=1}^{m}e^{i\phi_{[l]}}T_{lj}^{*}T_{lk},\quad j,k=1,2,\dots,m,\quad j\neq k.\end{split}$
@@
For a noncramped unitoid matrix, one can also write its SPD and GCF as above. However, neither SPD nor GCF is unique in this case. {eg} Consider the unitoid matrix $\begin{bmatrix}0&1\\ 1&0\end{bmatrix}$. For any $a>0$,
@@
In the case where $C$ is a real cramped matrix, sectoral decomposition, SPD, and GCF have their respective real counterparts. To see this, note that the numerical range of a real cramped matrix is symmetric about the real axis. Let
@@
which is a real normal matrix. In other words, $C$ is congruent to a real normal matrix via a real congruence. Since a real normal matrix $A$ can be decomposed [12, Theorem 2.5.8] into
@@
it follows that a real cramped matrix $C$ can be factorized as
@@
where $T$ is a nonsingular real matrix and $D$ is a real block orthogonal matrix with each block being either a scalar or a 2-by-2 matrix. We call this a real sectoral decomposition, which is nonunique in general. By performing the real polar decomposition and real QR decomposition of $T$, we arrive, respectively, at the real SPD and GCF, whose uniqueness follows from Theorem 3.1. To be specific, there exist unique real positive definite $P$, real upper triangular $R$ with positive diagonal elements, and real orthogonal $U$ and $W$ such that

### v2

@@
Abstract
In this paper, we define the phases of a complex sectorial matrix to be its canonical angles, which are uniquely determined from a sectorial decomposition of the matrix. Various properties of matrix phases are studied, including those of compressions, Schur complements, matrix products, and sums. In particular, by exploiting a notion known as the compound numerical range, we establish a majorization relation between the phases of the eigenvalues of $AB$ and the phases of $A$ and $B$. This is then applied to investigate the rank of $I+AB$ with phase information of $A$ and $B$, which plays a crucial role in feedback stability analysis. A pair of problems: banded sectorial matrix completion and decomposition is studied. The phases of the Kronecker and Hadamard products are also discussed.
@@ Abstract
phases , arguments , sectorial matrices , interlacing , majorization , rank robustness
2010 MSC:
15A18 , 15A23 , 15A42 , 15A60 , 15B48 , 15B57 , 15A83
@@ 1 Introduction
To be more specific, the phase $\phi$ takes values in a half open $2\pi$-interval, typically $0,2\pi)$ or $(-\pi,\pi]$. If $c=0$, the phase $\angle c$ is undefined.
@@
$|ab|$ $=|a||b|$ (multiplicativity) (1)
$|a+b|$ $\leq|a|+|b|$ (subadditivity) (2)
$\angle(ab)$ $=\angle a+\angle b\ \ \ \mbox{ mod $2\pi$}$ (3)
$\angle(a+b)$ $\in\mathrm{Co}\{\angle a,\angle b\}\ \ \ \mbox{if $|\angle a-\angle b|<\pi$},$ (4)
@@
It is well accepted that an $n\times n$ complex matrix $C$ has $n$ magnitudes, served by the $n$ singular values
@@
The singular values can be obtained from a singular value decomposition $C=USV^{*}$, where $U,V$ are unitary, and $S=\mathrm{diag}\left\{\sigma_{1}(C),\sigma_{2}(C),\dots,\sigma_{n}(C)\right\}$ [1, Theorem 2.6.3]. They can also be derived from the following maximin and minimax expressions [2, Corollary III.1.2][1, Theorem 4.2.6]:
@@
Note that log-majorization is stronger than weak-majorization, i.e., $x\prec_{\log}y$ implies $x\prec_{w}y$ [3, Chapter 5, A.2.b]. For more details on the theory of majorization, we refer to [3] and the references therein. Denote the vector of eigenvalues of $C$ by
@@
It is well known that the magnitudes of the eigenvalues are bounded by the singular values in the following manner [3, Theorem 9.E.1]:
@@
The singular values of $AB$ and $A+B$ are related to those of $A$ and $B$ through the following majorization type inequalities resembling (1) and (2), respectively, [3, Theorems 9.H.1 and 9.G.1.d]
$\sigma(AB)$ $\prec_{\log}\sigma(A)\odot\sigma(B),$ (5)
$\sigma(A+B)$ $\prec_{w}\sigma(A)+\sigma(B),$ (6)
@@
In contrast to the magnitudes of a complex matrix $C$, there does not exist a universally accepted definition of phases of $C$. What properties should the phases satisfy? In this paper, we advocate a definition of matrix phases and derive some of their properties as counterparts to those of the singular values.
An early attempt [4] defined the phases of $C$ as the phases of the eigenvalues of the unitary part of its polar decomposition, as motivated by the seeming generalization of the polar form of a scalar to the polar decomposition of a matrix. As in the scalar case, ambiguity arises when defining phases for a singular $C$. Hence, only nonsingular matrices are relevant. More precisely, let the left and right polar decompositions of a nonsingular matrix $C$ be given by $C=PU=UQ$, where $P$ and $Q$ are positive definite and $U$ is unitary [1, Theorem 7.3.1]. The authors of [4] proposed to define the phases of $C$ as
$\psi(C)=\angle\lambda(U).$ (7)
@@
In this paper, we shall revisit the matrix canonical angles introduced in [5] and propose to adopt them as the phases of a complex matrix. The phases defined in this way have the desired properties as shown in earlier studies and this paper.
@@ 2 Definition of phases
First, we review the concepts of numerical range, angular numerical range, and canonical angles of a matrix [6, Chapter 1]. The numerical range, also called the field of values, of a matrix $C\in\mathbb{C}^{n\times n}$ is defined as
$W(C)=\{x^{*}Cx:x\in\mathbb{C}^{n},\|x\|=1\},$
@@
If $0\notin W(C)$, then $W(C)$ is completely contained in an open half plane by its convexity. In this case, $C$ is said to be a sectorial matrix [7], also called sector matrix in [8]. The sectorial matrices have been widely studied in the literature [9, 10, 11].
Let $C$ be a sectorial matrix. It is well known [12, 13, 9] that $C$ is congruent to a unitary matrix. In particular, the unitary matrix can be chosen to be diagonal, i.e., there exist a nonsingular matrix $T$ and a diagonal unitary matrix $D$ such that
$C=T^{*}DT.$ (8)
The factorization (8) is called sectorial decomposition [11]. It is clear that $W^{\prime}(D)=W^{\prime}(C)$ and thus $D$ is also sectorial. Let $\delta(C)$ be the field angle of $C$, i.e., the angle subtended by the two supporting rays of $W(C)$ at the origin. We define the phases of $C$, denoted by $\phi_{1}(C),\phi_{2}(C),\dots,\phi_{n}(C)$, to be the phases of the eigenvalues (i.e., diagonal elements) of $D$, taking values in an interval $(\theta,\theta+\pi)$ for some $\theta\in[-\pi,\delta(C))$. The phases defined in this way coincide with the canonical angles of $C$ introduced in [5].
A sectorial decomposition for a sectorial matrix is not unique. Nevertheless, the diagonal unitary matrix $D$ is unique up to a permutation, as pointed out in [11]. As such, the phases of a sectorial matrix are uniquely defined. We adopt the convention of labeling the phases in such a way that
@@
Define $\phi(C)=[\phi_{1}(C)\ \ \phi_{2}(C)\ \ \dots\ \ \phi_{n}(C)]$. The uniqueness of the phases also follows directly from the maximin and minimax expressions [12, Lemma 8]:
$\begin{split}\phi_{i}(C)&=\max_{\mathcal{M}:\mathrm{dim}\mathcal{M}=i}\min_{x\in\mathcal{M},\|x\|=1}\angle x^{*}Cx\\ &=\min_{\mathcal{N}:\mathrm{dim}\mathcal{N}=n-i+1}\max_{x\in\mathcal{N},\|x\|=1}\angle x^{*}Cx.\end{split}$ (9)
@@
from which we can observe that $\overline{\phi}(C)$ and $\underline{\phi}(C)$ enjoy good geometric interpretations. Consider the two supporting rays of $W(C)$. Since $0\notin W(C)$, both supporting rays can be determined uniquely. Figure 1 illustrates an example of $W(C)$ and its supporting rays. The two angles from the positive real axis to the supporting rays are $\overline{\phi}(C)$ and $\underline{\phi}(C)$ respectively. The other phases of $C$ lie in between.
Figure 1: Geometric interpretations of $\overline{\phi}(C)$ and $\underline{\phi}(C)$.
Given matrix $C$, we can check whether it is sectorial or not by plotting its numerical range. From the plot of numerical range, we can also determine a $\pi$-interval $(\theta,\theta+\pi)$ in which the phases take values. How to efficiently compute $\phi(C)$ is an important issue. The following observation provides some insights along this direction. Suppose $C$ is sectorial. Then it admits a sectorial decomposition $C=T^{*}DT$ and thus
$C^{-1}C^{*}=T^{-1}D^{-1}T^{-*}T^{*}D^{*}T=T^{-1}D^{-2}T,$
indicating that $C^{-1}C^{*}$ is similar to a diagonal unitary matrix. Hence, we can first compute $\angle\lambda(C^{-1}C^{*})$, taking values in $(-2\theta\!-\!2\pi,-2\theta)$, and let $\phi(C)=-\frac{1}{2}\angle\lambda(C^{-1}C^{*})$. This gives one possible way to compute $\phi(C)$. We are currently exploring other methods, hopefully of lower complexity, for determination of the interval in which matrix phases take values.
It is worth noting that the class of unitoids was introduced in [14] to consist of matrices that are congruent to unitary matrices. Clearly, sectorial matrices constitute a special type of unitoid matrices. A nonsectorial unitoid matrix admits a factorization of the form (8). However, in this case, the eigenvalues of $D$ lie on an arc of the unit circle of length no less than $\pi$. In this paper, we do not define the phases of such matrices.
Example 2.1
@@
is not sectorial. We do not define the phases of this matrix, since in this endeavor there would clearly be an ambiguity in deciding whether the phases should be taken as $2\pi/3,0,-2\pi/3$ or $4\pi/3,2\pi/3,0$.
@@
    The phases of a sectorial normal matrix are the phases of its eigenvalues.
@@
    The phases of $C$ belong to $(-\pi/2,\pi/2)$ if and only if $C$ has a positive definite Hermitian part, i.e., $(C+C^{*})/2>0$. Such matrices are called strictly accretive matrices [15] [16, p. 279].
@@
$\psi\left(\mbox{$\left[\begin{array}[]{cc}\cos\alpha&0\\ 0&\sin\alpha\end{array}\right]$}\mbox{$\left[\begin{array}[]{cc}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{array}\right]$}\mbox{$\left[\begin{array}[]{cc}\cos\alpha&0\\ 0&\sin\alpha\end{array}\right]$}\right)\\ =\begin{bmatrix}\arccos\displaystyle\frac{\cos\theta}{\sqrt{1-\sin^{2}\theta\cos^{2}2\alpha}}&\;-\arccos\displaystyle\frac{\cos\theta}{\sqrt{1-\sin^{2}\theta\cos^{2}2\alpha}}\end{bmatrix}$, (10)
@@
which is negative for some $\theta$ close to $\pi/2$ and $\alpha$ away from $\pi/4$, such as $\theta=\pi/4$ and $\alpha=\pi/12$. That means the matrix is not strictly accretive for these values of $\theta$ and $\alpha$.
The following lemma shows that the phases of a sectorial matrix provide a bound on the phases of its eigenvalues, just as the magnitudes of a matrix provide a bound on the magnitudes of its eigenvalues. It has been proved implicitly in [12, Lemma 9] and also shown in [5, Theorem 1].
Lemma 2.3
Let $C$ be sectorial with phases in $(\theta,\theta+\pi)$, where $\theta\in[-\pi,\delta(C))$. Then the phases of the eigenvalues of $C$ can be chosen so that
@@
in Example 2.2, where $\alpha,\theta\in(0,\frac{\pi}{2})$. One can see from the discussion therein that $\psi(C)\prec\phi(C)$. In fact, this is generally true for any sectorial matrix $C$ with phases in $(\theta,\theta+\pi)$, if $\psi(C)$ takes values in the same interval, which can be inferred from [12, Lemma 9].
3 Sectorial matrix decompositions
As discussed in the previous section, a sectorial matrix $C$ admits a sectorial decomposition as in (8), which is not unique. If the unitary matrix in (8) is not restricted to be diagonal, one gains more flexibility. In fact, letting $T=VP$ be the unique right polar decomposition and $U=V^{*}DV$ leads to a new decomposition
@@ 3 Sectorial matrix decompositions
where $P$ is positive definite and $U$ is unitary. This is called the symmetric polar decomposition (SPD). Meanwhile, letting $T=QR$ be the unique QR factorization [1, Theorem 2.1.14] with $Q$ being unitary and $R$ being upper triangular with positive diagonal elements and letting $W=Q^{*}DQ$ leads to another decomposition
@@
Theorem 3.1
The SPD and GCF of a sectorial matrix $C$ are unique.
Proof 1
@@
Since the phases of a sectorial matrix are unique, we can find unitary matrices $W_{1}$ and $W_{2}$ such that $U_{1}=W^{*}_{1}DW_{1}$ and $U_{2}=W^{*}_{2}DW_{2}$, where
@@
$\begin{split}e^{i\phi_{[k]}}I_{n_{k}}&=\sum_{l=1}^{m}e^{i\phi_{[l]}}T^{*}_{lk}T_{lk},\quad k=1,2,\dots,m\quad\text{and}\\ 0&=\sum_{l=1}^{m}e^{i\phi_{[l]}}T_{lj}^{*}T_{lk},\quad j,k=1,2,\dots,m,\quad j\neq k.\end{split}$ (12)
@@
For a nonsectorial unitoid matrix, one can also write its SPD and GCF as above. However, neither SPD nor GCF is unique in this case.
Example 3.2
Consider the unitoid matrix $\begin{bmatrix}0&1\\ 1&0\end{bmatrix}$. For any $a>0$,
@@
In the case where $C$ is a real sectorial matrix, sectorial decomposition, SPD, and GCF have their respective real counterparts. To see this, note that the numerical range of a real sectorial matrix is symmetric about the real axis. Let
@@
which is a real normal matrix. In other words, $C$ is congruent to a real normal matrix via a real congruence. Since a real normal matrix $A$ can be decomposed [1, Theorem 2.5.8] into
@@
it follows that a real sectorial matrix $C$ can be factorized as
@@
where $T$ is a nonsingular real matrix and $D$ is a real block-diagonal orthogonal matrix with each block being either a scalar or a 2-by-2 matrix. We call this a real sectorial decomposition, which is nonunique in general. By performing the real polar decomposition and real QR decomposition of $T$, we arrive, respectively, at the real SPD and GCF, whose uniqueness follows from Theorem 3.1. To be specific, there exist unique real positive definite $P$, real upper triangular $R$ with positive diagonal elements, and real orthogonal $U$ and $W$ such that